OpenStudy (anonymous):

Find an equation for the line tangent to y=-2-4x^2 at the point (-3,-38) i know how to set up the problem with f(x+h)-f(x)/h but i think i keep messing up a step while simplifying because i've gotten the answer wrong twice now

OpenStudy (isaiah.feynman):

Well is it a must f(x+h)=f(x)/h be used?

OpenStudy (anonymous):

that's the only way that i know how to solve it and its f(x+h)-f(x)/h

OpenStudy (anonymous):

if you know another way to solve it that would be fine

OpenStudy (isaiah.feynman):

I know another way, but you'd be confused. Lets use this way. Show me what you have done.

OpenStudy (anonymous):

i simplified it down to -2-4(h^2-6h+9)+38 and i don't know what to do next

OpenStudy (anonymous):

all of that over h too

OpenStudy (isaiah.feynman):

Well I continued from where you stopped, got a strange look! Let's start from the beginning.

OpenStudy (isaiah.feynman):

\[\lim_{h \rightarrow 0}\frac{ (-2-4(x+h)^{2})-(-2-4x^{2}) }{ h }\]

OpenStudy (anonymous):

ok i see where i messed up i didn't include the -2 in front of the f(x+h)

OpenStudy (anonymous):

i just plugged x into 4x^2

OpenStudy (isaiah.feynman):

Okay, try solving it now.

OpenStudy (anonymous):

should i distribute the 4 before i take (x+)^2

OpenStudy (anonymous):

(x+h) i mean

OpenStudy (isaiah.feynman):

No. Expand the bracket first.

OpenStudy (anonymous):

so -6(h^2-6h+9)-38 all over h?

OpenStudy (anonymous):

is that what you got to so far

OpenStudy (isaiah.feynman):

\[\lim_{h \rightarrow 0} \frac{ (-2-4(x^{2}+2xh+h^{2}))-(-2-4x ^{2}) }{ h }\]

OpenStudy (anonymous):

shouldn't you plug the -3 in as x before squaring it

OpenStudy (anonymous):

because it gives the point (-3,-38)

OpenStudy (isaiah.feynman):

Not now. Plug in 3 when you have evaluated that limit.

OpenStudy (anonymous):

so -4h-2x is what i got

OpenStudy (isaiah.feynman):

\[\lim_{h \rightarrow 0} \frac{ -2-4x^{2}-8xh-4h^2+2+4x^{2} }{ h }\]

OpenStudy (anonymous):

so everything but the -8xh and -4h^2 cancel out

OpenStudy (anonymous):

over h

OpenStudy (isaiah.feynman):

Simplifies to \[\lim_{h \rightarrow 0}\frac{ -8xh-4h^{2} }{ h }\]

OpenStudy (anonymous):

then i took out -4h(h-2x) over h

OpenStudy (isaiah.feynman):

Factoring out h from the top and cancelling gives..\[\lim_{h \rightarrow 0} (-8x-4h)=-8x\]

OpenStudy (isaiah.feynman):

Now substitute -3 for x like you've always wanted to.

OpenStudy (anonymous):

h=0?

OpenStudy (isaiah.feynman):

At that point the function is continuous so direct substitution can be used.

OpenStudy (anonymous):

what do i substitute?

OpenStudy (anonymous):

-3 into the original function?

OpenStudy (isaiah.feynman):

No. Into -8x. So you get -8(-3).

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

is 24 the slope then?

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

would that make the final equation y=24x+34

OpenStudy (isaiah.feynman):

No. Remember this? \[y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

ya i put it into y=mx+b form from point slope form

OpenStudy (isaiah.feynman):

y1=-38,x1=-3, m=24. Plug them in, re arrange and put it in slope intercept form (y=mx+b)

OpenStudy (anonymous):

i still got the same final equation

OpenStudy (isaiah.feynman):

Hopefully that's correct. That's the end of the problem.

OpenStudy (anonymous):

yep i just put it in thanks for the help

OpenStudy (isaiah.feynman):

Welcome.